OpenStudy (queenbeedee):

PLEASE HELP ME !! The point (1, −1) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Explain ?

OpenStudy (anonymous):

sin=√2/2 cos=-√2/2 tan=-1

OpenStudy (queenbeedee):

how did you get those answers ? @Lovelarap

OpenStudy (anonymous):